[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 22, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Jul 22 20:49:02 UTC 2017


HSU Shift Log for Sat Jul 22 2017 16:48:54 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Note the microwave data is getting into NAWIPS faster  
than NRL/FNMOC at times. FSSE arrived very late for Greg at 12z-- after  
1400Z, too late for the fcst package. With the formation of 10-E, we have  
opened 11 EPac invests and gotten 10 TCs (out of only 13 five-day  
systems)!!!
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170722/538bcb4f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list